1. The selling price is calculated at a mark-up of 50% on variable cost.

2. New products introduced must contribute 15% net profit after tax.


He supplied you with the following information:

   Cost per unit   Other information

Leather   R246.00   

Silk material   R350.00   

Labour   3 hours @ R185.00 per hour   

Overheads   R155.00 per labour hour   

Incremental fixed cost      R625,000.00

Tax rate      28%


The sales manager estimated the market size to be 2,000 handbags for the year.

Calculate and comment if the handbag will comply with the company’s policy regarding the introduction of new products.


1
Expert's answer
2021-07-01T05:05:01-0400

variable cost:

"246+350+3*185+3*155=1616"

Profit per unit:

"1616*0.5=808"

Profit before tax:

"2000*808-625000=991000"

Profit after tax:

"991000*(1-0.28)=713520"

Sales:

"(1616+808)*2000=4848000"

Profitability:

"\frac{713520}{4848000}="0.1472=14.72%<15%

handbag will not comply with the company’s policy